# Causal Data Fusion for Panel Data without a Pre-Intervention Period

### v3

@@ 1 Introduction
The difference-in-differences framework (Card,, 1990; Angrist and Pischke,, 2008) and its major extension, the synthetic control framework (Abadie and Gardeazabal,, 2003; Abadie et al.,, 2010, 2015), are well-known panel-data methods commonly used in such settings. These methods require access to pre-intervention data from the target unit to match it with units (or combinations of units) from the control group. However, in some applications, such pre-intervention data may not be available. For example, in the case of a sudden shock or emergency, immediate action may be taken by agencies, leaving no data from the unit during the emergency before the action was implemented. During the COVID-19 pandemic, interventions were implemented almost instantaneously to curb the spread of the virus. When COVID-19 vaccines first became available in Massachusetts, doses were accessible primarily to English-fluent people who could use online appointment scheduling sites and who could drive to mass vaccination centers. For many low-income residents who lacked access to information in their preferred language, to a computer, or to a car, COVID-19 vaccines remained out of reach. Efforts by a community-based organization in the city of Chelsea became exemplary for improving vaccine access for their community members. As the organization began advocating for vaccine access and mounting a campaign to provide vaccine information for their community members (which is the intervention in the system) as soon as vaccines became available to the general public in January 2021, these circumstances left researchers with no pre-intervention data from the target unit to analyze the impact of the interventions. This lack of pre-intervention data hampers the ability to draw robust conclusions about the effectiveness of the interventions and to inform future policy decisions.
@@
The use of data fusion for identification and estimation of causal parameters when data from a single domain are insufficient for effective evaluation has gained significant attention in recent years (Dahabreh et al.,, 2020; Degtiar and Rose,, 2023; Colnet et al.,, 2024). However, to the best of our knowledge, we are the first to study this approach in the challenging setting of panel data with a single target unit where latent confounding variables may be present.
@@ 2 Problem Description
We refer to unit $i=1$ as the target unit. An example of this setting is the study of the effect of community organizations on the COVID-19 vaccination rate of a sub-population in cities, as mentioned in Section 1. Here, the target city had a strong community organization with a large team of community health workers and the capacity to provide wraparound services addressing fundamental social determinants of health from the beginning of the vaccine rollout, whereas the remaining cities under consideration did not have such community organizations.
We adopt the potential outcomes framework (Rubin,, 1974) and denote the potential outcome of $Y_{i,s}$, had the treatment variable been set to $A_{i,s}=a$ for all $s$ (possibly contrary to fact), by $Y_{i,s}^{(a)}$, where $a\in\{0,1\}$. Our parameter of interest is the effect of the treatment on the treated (ETT) averaged over the $S$ treatment time periods:
@@
We assume the consistency assumption (Hernán and Robins,, 2020), formally stated as follows.
@@
By consistency, $\psi_{0}$ can be written as $\psi_{0}=\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}[Y_{1,s}-Y_{1,s}^{(0)}]$. Hence, the challenge in evaluating the causal effect lies in assessing the potential outcome mean of the target unit under control. Similar settings have been considered in the literature on causal inference within the difference-in-differences framework, the synthetic control framework, and their variations (Arkhangelsky and Imbens,, 2024). However, in those approaches it is assumed that the target unit is also observed before the implementation of the treatment, which is not the case in our setting. Therefore, those approaches cannot be directly applied here.
In our setting of interest, we also have access to data from a secondary domain for the same units. We refer to the primary domain, in which we aim to evaluate the causal parameter, as the target domain, and to the secondary domain as the reference domain. In the reference domain, units are observed over $T$ time periods. We denote the observed covariates of the units in this domain by $Z_{i}\in\mathbb{R}^{d_{r}}$, $1\leq i\leq J+1$. Note that the covariates observed in the target and reference domains need not be the same, and $d_{r}$ is not necessarily equal to $d_{t}$. We also observe an outcome variable $F_{i,t}\in\mathbb{R}$ for $1\leq i\leq J+1$ and $1\leq t\leq T$ in the reference domain.
@@ 3.1 Linear Equi-Confounding
Our first proposed approach extends the ideas of negative controls and the difference-in-differences causal inference framework to our data fusion setup. A similar extension appeared in (Ghassami et al.,, 2022) for evaluating the long-term causal effect of treatments via data fusion.
@@
The left-hand side of Equation (2) involves variables in the target domain, while the right-hand side involves variables in the reference domain. Hence, Assumption 2 specifies the connection between the outcomes in the two domains and clarifies how information in the reference domain can be transferred to the target domain. The quantity $\mathbb{E}[Y_{1}^{(0)}]-\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[Y_{i}^{(0)}]$ measures the confounding effect on the potential outcome $Y^{(0)}_{1}$. Therefore, Assumption 2 requires that this effect for $Y^{(0)}_{1}$ and $F_{1}$ be equal; hence the name equi-confounding. The assumption is inspired by the additive equi-confounding bias assumption in the identification approach using negative outcome controls (Lipsitch et al.,, 2010; Tchetgen Tchetgen,, 2014; Sofer et al.,, 2016), and is similar in spirit to the parallel trends assumption in the difference-in-differences framework (Card,, 1990; Angrist and Pischke,, 2008).
@@
$Y_{i,s}=\varrho_{s}+f(G_{i},X_{i})+\alpha_{i,s}A_{i,s}+\varepsilon_{i,s},\quad F_{i,t}=\rho_{t}+g(G_{i},Z_{i})+\epsilon_{i,t},$
with $X_{i}\overset{i.i.d.}{\sim}P_{X}$ and $Z_{i}\overset{i.i.d.}{\sim}P_{Z}$, for some distributions $P_{X}$ and $P_{Z}$, and $\mathbb{E}[f(1,X_{i})-f(0,X_{i})]=\mathbb{E}[g(1,Z_{i})-g(0,Z_{i})]$, where $\varepsilon_{i,s}$ and $\epsilon_{i,t}$ are mean-zero independent noise variables, and $G_{i}$ is a target-unit indicator variable equal to $1$ if $i=1$ and 0 otherwise.
@@ 3.2 Logarithmic Equi-Confounding
Similar to Equation (2), in Equation (3) the left-hand side involves variables in the target domain and the right-hand side involves variables in the reference domain. Hence, Assumption 3 specifies the connection between outcomes in the two domains. This assumption can be viewed as a modification of Assumption 2, in which the change in the mean is replaced by the change in the logarithm of the mean; hence the name logarithmic equi-confounding. A similar observation has been made in (Wooldridge,, 2023). More specifically, Equation (3) can be written as
@@
  3. 3.
    There exists a function $C$ with $C(J)\downarrow 0$ as $J\to\infty$, such that
    $\left|\mathrm{cov}\!\left(\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}},\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}\right)\right|\;\leq\;C(J).$
Part (2) of Assumption 4 rules out a fully connected dependence structure among control units. This assumption is trivially satisfied if the outcomes for the control units in the reference domain are pairwise independent. Regarding Part (3), note that this assumption does not require the outcome variables in the target and reference domains to be (approximately) uncorrelated. Instead, it requires that the sample mean $1/J\sum_{i=2}^{J+1}Y_{i}$ be approximately uncorrelated with the ratio quantity $F_{1}/(1/J\sum_{i=2}^{J+1}F_{i})$, which is a weaker requirement, as the ratio can remain approximately unchanged even when the individual variables vary.
@@
where $\Delta_{f}:=L_{f}-l_{f}$. In particular, as $J\to\infty$ with $\tau(J)\to 0$ and $C(J)\to 0$, the bias vanishes at the rate $O(J^{-1/2})$ up to $\tau(J)$ and $C(J)$.
This result demonstrates that, under the stated assumptions on the data-generating process, the bound on the bias of our proposed estimator vanishes as the number of control units increases. The bound in Theorem 3 can be further improved if we also bound the cross-covariance between the target and control units:
@@ 4 Synthetic Control Data Fusion
The equi-confounding assumptions introduced in the previous section require a relation between the average of the outcomes for the control units and the outcomes of the target unit in the reference and target domains. In this averaging, all units are weighted equally, with weights $1/J$. In many applications, certain control units are more similar to the target unit and, hence, may contain more information regarding the counterfactual outcomes of the target unit. Therefore, a natural extension of the ideas in the previous section is to consider different weights for different control units. This is the main idea of the synthetic control method (Abadie and Gardeazabal,, 2003; Abadie et al.,, 2010, 2015). Our goal is to formalize this idea in our data fusion setting.
We consider weight $w_{i}$ for the $i$-th control unit, $2\leq i\leq J+1$, and define the $J$-dimensional weight vector $w=[w_{2}~\cdots~w_{J+1}]^{\top}$, such that the weights satisfy $w_{i}\geq 0$ for $i=2,\dots,J+1$, and $\sum_{i=2}^{J+1}w_{i}=1$. Our goal is to choose weights such that the restrictions below are (approximately) satisfied:
  • The weighted sum of the outcomes of the control units in the reference domain at time $t$ matches that of the target unit, for $1\leq t\leq T$, i.e.,
    $F_{1,t}=\sum_{i=2}^{J+1}w_{i}F_{i,t},\qquad 1\leq t\leq T.$ (4)
  • The weighted sum of the observed covariates of the control units in the reference domain matches that of the target unit, i.e.,
    $Z_{1}=\sum_{i=2}^{J+1}w_{i}Z_{i}.$ (5)
  • The weighted sum of the observed covariates of the control units in the target domain matches that of the target unit, i.e.,
    $X_{1}=\sum_{i=2}^{J+1}w_{i}X_{i}.$ (6)
The intuition for imposing the above restrictions is as follows. Suppose the matching requirements are satisfied and the outcomes of the units in the reference and target domains are generated by mechanisms that differ primarily in how they are affected by the observed covariates. In this case, if we match on the observed covariates and the outcomes over a long period in the reference domain, then we should also be matching on any unobserved covariates that affect outcomes in both domains. For instance, in the vaccination-rate application, with the Hispanic and Black sub-populations as the target and reference domains, respectively, factors such as overall attitudes toward vaccines, socioeconomic status, and access to vaccines are unobserved covariates that may similarly affect COVID-19 vaccination rates across both sub-populations. Therefore, if we match on the vaccination rate from one subpopulation for an extended period, as well as on all observed covariates relevant to COVID-19 vaccination, then there is statistical justification that we are also matching on unobserved covariates.
Based on the above argument, a weighted sum of the potential outcomes with weight vector $w$ that (approximately) satisfies restrictions (4)–(6) will match the potential outcome of the target unit to a good approximation. That is,
@@
In restrictions (4)–(6), we will usually have more equations than unknowns, and, hence, these restrictions cannot be exactly satisfied. Therefore, we instead minimize the mismatch in each restriction. Formally, for $i=1,\dots,J+1$, let $\bar{F}_{i}:=[F_{i,1}~\cdots~F_{i,T}]^{\top}$, and $K_{i}:=[\bar{F}_{i}^{\top}~Z_{i}^{\top}~X_{i}^{\top}]^{\top}$. A naive way to obtain the weight vector $w$ is
@@
In the approach above, in line with the original synthetic control method, we require the weights $w$ to be nonnegative and to sum to one. The purpose of this restriction is interpretability and avoiding overfitting. Yet, similar to the standard synthetic control method, we can also prevent overfitting by regularization (Doudchenko and Imbens,, 2016).
@@
Here, $\alpha_{i,s}$ is the causal effect on unit $i$ in time period $s$. The vectors $\mu_{i}\in\mathbb{R}^{d_{u}}$ are unit-specific latent factors that affect both domains; $\tilde{\mu}_{i}^{Y}\in\mathbb{R}^{d^{Y}_{\tilde{u}}}$ and $\tilde{\mu}_{i}^{F}\in\mathbb{R}^{d^{F}_{\tilde{u}}}$ are unit-specific latent factors that affect only the target and reference domains, respectively. Loadings are deterministically bounded in $\ell_{2}$-norm: $\sup_{t\leq T}\|\theta_{t}\|_{2}\leq\bar{\theta}$, $\sup_{t\leq T}\|\tilde{\theta}_{t}\|_{2}\leq\bar{\tilde{\theta}}$, $\sup_{s\leq S}\|\vartheta_{s}\|_{2}\leq\bar{\vartheta}$, and $\sup_{s\leq S}\|\tilde{\vartheta}_{s}\|_{2}\leq\bar{\tilde{\vartheta}}$. The matrix $\sum_{t=1}^{T}\theta_{t}\theta_{t}^{\top}$ has minimum eigenvalue $\lambda_{\min}\geq T\,\underline{\xi}>0$. The shocks satisfy $\mathbb{E}[\epsilon_{i,t}]=\mathbb{E}[\varepsilon_{i,s}]=0$, and $\epsilon_{i,t}$ are i.i.d. sub-Gaussian with variance proxy $\bar{\sigma}^{2}$. The collections $\{\tilde{\mu}^{Y}_{i}\}_{i\leq J+1}$ and $\{\tilde{\mu}^{F}_{i}\}_{i\leq J+1}$ are i.i.d., mean-zero, sub-Gaussian with variance proxy $\tau^{2}$. Let $\mathcal{G}:=\sigma\!\big(\{X_{i},Z_{i},\mu_{i},\tilde{\mu}^{F}_{i}\}_{i\leq J+1},\{\epsilon_{i,t}\}_{i\leq J+1,\,1\leq t\leq T}\big)$ denote the $\sigma$-field generated by all reference-domain objects used to construct the synthetic-control weights $w=(w_{2},\ldots,w_{J+1})$, which are assumed $\mathcal{G}$-measurable, nonnegative, and satisfy $\sum_{i=2}^{J+1}w_{i}=1$. We assume that $\{\varepsilon_{i,s}\}_{i,s}$ and $\{\tilde{\mu}^{Y}_{i}\}_{i}$ are independent of $\mathcal{G}$.
@@
$\big|\mathbb{E}[\hat{\psi}^{sc}]-\psi_{0}\big|\;\leq\;\sqrt{2}\,\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,\bar{\sigma}\sqrt{\frac{\log(2J)}{T}}\;+\;2\,\frac{\bar{\vartheta}\,\bar{\theta}\,\bar{\tilde{\theta}}}{\underline{\xi}}\,\tau\,\sqrt{\log(2J)}.$
Theorem 4 shows that, in the absence of reference-domain–only latent factors ($\tilde{\mu}^{F}_{i}$), the bound on the bias of our proposed estimator vanishes as the number of time periods in the reference domain increases. By contrast, the presence of target-domain–only latent factors ($\tilde{\mu}^{Y}_{i}$) does not prevent convergence. We also see that the bound becomes wider as the number of donor units increases, due to the higher chance of overfitting, i.e., matching by chance. A main ingredient in the proof is a maximal inequality to bound a linear combination of weighted error terms. This step also appears in Abadie et al., (2010) in the standard synthetic control setting. However, in contrast to that work, we employ a sub-Gaussian approach (see Assumption 5), leading to a clearer and more interpretable derivation without additional moment assumptions on auxiliary variables used in Abadie et al., (2010). There are also challenges unique to the data fusion setting that do not arise in the standard synthetic control setting: (i) we must handle target-domain–only and reference-domain–only latent factors; and (ii) we must account for two sets of factor loadings for the shared latent factors $\mu_{i}$ across the two domains, which necessitates bounds for both sets. As seen in Theorem 4, the upper bounds on both sets of loadings appear in the bias bound.

### v4

@@ 1 Introduction
The difference-in-differences framework (Card, 1990; Angrist and Pischke, 2008) and its major extension, the synthetic control framework (Abadie and Gardeazabal, 2003; Abadie et al., 2010, 2015), are well-known panel-data methods commonly used in such settings. These methods require access to pre-intervention data from the target unit to match it with units (or combinations of units) from the control group. However, in some applications, such pre-intervention data may not be available. For example, in the case of a sudden shock or emergency, immediate action may be taken by agencies, leaving no data from the unit during the emergency before the action was implemented. During the COVID-19 pandemic, interventions were implemented almost instantaneously to curb the spread of the virus. When COVID-19 vaccines first became available in Massachusetts, doses were accessible primarily to English-fluent people who could use online appointment scheduling sites and who could drive to mass vaccination centers. For many low-income residents who lacked access to information in their preferred language, to a computer, or to a car, COVID-19 vaccines remained out of reach. Efforts by a community-based organization in the city of Chelsea became exemplary for improving vaccine access for their community members. As the organization began advocating for vaccine access and mounting a campaign to provide vaccine information for their community members (which is the intervention in the system) as soon as vaccines became available to the general public in January 2021, these circumstances left researchers with no pre-intervention data from the target unit to analyze the impact of the interventions. This lack of pre-intervention data hampers the ability to draw robust conclusions about the effectiveness of the interventions and to inform future policy decisions.
@@
The use of data fusion for identification and estimation of causal parameters when data from a single domain are insufficient for effective evaluation has gained significant attention in recent years (Dahabreh et al., 2020; Degtiar and Rose, 2023; Colnet et al., 2024). However, to the best of our knowledge, we are the first to study this approach in the challenging setting of panel data with a single target unit where latent confounding variables may be present.
@@ 2 Problem Description
We refer to unit $i=1$ as the target unit. An example of this setting is the study of the effect of community organizations on the COVID-19 vaccination rate of a sub-population in cities, as mentioned in Section 1. Here, the target city (Chelsea) had a strong community organization with a large team of community health workers and the capacity to provide wraparound services addressing fundamental social determinants of health from the beginning of the vaccine rollout, whereas the remaining cities under consideration did not have such community organizations.
We adopt the potential outcomes framework (Rubin, 1974) and denote the potential outcome of $Y_{i,s}$, had the treatment variable been set to $A_{i,s}=a$ for all $s$ (possibly contrary to fact), by $Y_{i,s}^{(a)}$, where $a\in\{0,1\}$. Our parameter of interest is the effect of the treatment on the treated (ETT) averaged over the $S$ treatment time periods:
@@
We assume the consistency assumption (Hernán and Robins, 2020), formally stated as follows.
@@
By consistency, $\psi_{0}$ can be written as $\psi_{0}=\frac{1}{S}\sum_{s=1}^{S}\mathbb{E}[Y_{1,s}-Y_{1,s}^{(0)}]$. Hence, the challenge in evaluating the causal effect lies in assessing the potential outcome mean of the target unit under control. Similar settings have been considered in the literature on causal inference within the difference-in-differences framework, the synthetic control framework, and their variations (Arkhangelsky and Imbens, 2024). However, in those approaches it is assumed that the target unit is also observed before the implementation of the treatment, which is not the case in our setting. Therefore, those approaches cannot be directly applied here.
In our setting, we also have access to data from a secondary domain for the same units. We refer to the primary domain, in which we aim to evaluate the causal parameter, as the target domain, and to the secondary domain as the reference domain. In our application, the target domain consists of the COVID-19 vaccination rates of the Hispanic sub-population in Massachusetts cities, while the reference domain is the vaccination rates of the Black sub-population in those same cities. In the reference domain, units are observed over $T$ time periods. We denote the observed covariates of the units in this domain by $Z_{i}\in\mathbb{R}^{d_{r}}$, $1\leq i\leq J+1$. Note that the covariates observed in the target and reference domains need not be the same, and $d_{r}$ is not necessarily equal to $d_{t}$. We also observe an outcome variable $F_{i,t}\in\mathbb{R}$ for $1\leq i\leq J+1$ and $1\leq t\leq T$ in the reference domain.
@@ 3.1 Linear Equi-Confounding
Our first proposed approach extends the ideas of negative controls and the difference-in-differences causal inference framework to our data fusion setup. A similar extension appeared in (Ghassami et al., 2022) for evaluating the long-term causal effect of treatments via data fusion.
@@
The left-hand side of Equation (2) involves variables in the target domain, while the right-hand side involves variables in the reference domain. Hence, Assumption 2 specifies the connection between the outcomes in the two domains and clarifies how information in the reference domain can be transferred to the target domain. The quantity $\mathbb{E}[Y_{1}^{(0)}]-\frac{1}{J}\sum_{i=2}^{J+1}\mathbb{E}[Y_{i}^{(0)}]$ measures the confounding effect on the potential outcome $Y^{(0)}_{1}$ (in our application, this could be, e.g., the effect of regional socioeconomic barriers or shared healthcare infrastructure). Therefore, Assumption 2 requires that this effect for $Y^{(0)}_{1}$ and $F_{1}$ be equal; hence the name equi-confounding. The assumption is inspired by the additive equi-confounding bias assumption in the identification approach using negative outcome controls (Lipsitch et al., 2010; Tchetgen Tchetgen, 2014; Sofer et al., 2016), and is similar in spirit to the parallel trends assumption in the difference-in-differences framework (Card, 1990; Angrist and Pischke, 2008).
@@
$Y_{i,s}=\varrho_{s}+\theta G_{i}+\varphi^{\top}X_{i}+\alpha_{i,s}A_{i,s}+\varepsilon_{i,s},\quad F_{i,t}=\rho_{t}+\theta G_{i}+\phi^{\top}Z_{i}+\epsilon_{i,t},$
with $X_{i}\overset{i.i.d.}{\sim}P_{X}$ and $Z_{i}\overset{i.i.d.}{\sim}P_{Z}$, for some distributions $P_{X}$ and $P_{Z}$, where $\varepsilon_{i,s}$ and $\epsilon_{i,t}$ are mean-zero independent noise variables, and $G_{i}$ is a target-unit indicator variable equal to $1$ if $i=1$ and 0 otherwise.
@@ 3.2 Logarithmic Equi-Confounding
Similar to Equation (2), in Equation (3) the left-hand side involves variables in the target domain and the right-hand side involves variables in the reference domain. Hence, Assumption 3 specifies the connection between outcomes in the two domains. This assumption can be viewed as a modification of Assumption 2, in which the change in the mean is replaced by the change in the logarithm of the mean; hence the name logarithmic equi-confounding. A similar observation has been made in (Wooldridge, 2023). More specifically, Equation (3) can be written as
@@
Part (2) of Assumption 4 rules out a fully connected dependence structure among control units. This assumption is trivially satisfied if the outcomes for the control units in the reference domain are pairwise independent.
Let $C(J)$, denote an upper bound for the quantity
$\left|\mathrm{cov}\!\left(\frac{F_{1}}{\frac{1}{J}\sum_{i=2}^{J+1}F_{i}},\frac{1}{J}\sum_{i=2}^{J+1}Y_{i}\right)\right|.$
@@
where $\Delta_{f}:=L_{f}-l_{f}$.
Theorem 3 demonstrates that as $J\to\infty$ the bias will be controlled by $C(J)$. Note that often it is expected that $C(J)$ also vanishes as $J\to\infty$. In this case, the bound on the bias of our proposed estimator vanishes as the number of control units increases. Note that this property does not require the outcome variables in the target and reference domains to be (approximately) uncorrelated. Instead, it requires that the sample mean $1/J\sum_{i=2}^{J+1}Y_{i}$ be approximately uncorrelated with the ratio quantity $F_{1}/(1/J\sum_{i=2}^{J+1}F_{i})$, which is a weaker requirement, as the ratio can remain approximately unchanged even when the individual variables vary.
The bound in Theorem 3 can be further improved if we also bound the cross-covariance between the target and control units:
@@ 4 Synthetic Control Data Fusion
The equi-confounding assumptions introduced in the previous section require a relation between the average of the outcomes for the control units and the outcomes of the target unit in the reference and target domains. In this averaging, all units are weighted equally, with weights $1/J$. In many applications, certain control units are more similar to the target unit and, hence, may contain more information regarding the counterfactual outcomes of the target unit. Therefore, a natural extension of the ideas in the previous section is to consider different weights for different control units. This is the main idea of the synthetic control method (Abadie and Gardeazabal, 2003; Abadie et al., 2010, 2015). Our goal is to formalize this idea in our data fusion setting.
We consider weights $w_{i}$ for the $i$th control unit, $2\leq i\leq J+1$, and define $w=[w_{2}~\cdots~w_{J+1}]^{\top}$, such that $w_{i}\geq 0$ for $i=2,\ldots,J+1$ and $\sum_{i=2}^{J+1}w_{i}=1$. Let $\bar{F}_{i}:=[F_{i,1}~\cdots~F_{i,T}]^{\top}\in\mathbb{R}^{T}$.
Our goal is to choose weights such that the following approximate matching requirements hold: there exists a constant $c\geq 0$ such that, with probability one,
$\left\|\bar{F}_{1}-\sum_{i=2}^{J+1}w_{i}\bar{F}_{i}\right\|_{2}^{2}$ $\leq Tc^{2},$ (4)
$\left\|Z_{1}-\sum_{i=2}^{J+1}w_{i}Z_{i}\right\|_{2}^{2}$ $\leq d_{r}c^{2},$ (5)
$\left\|X_{1}-\sum_{i=2}^{J+1}w_{i}X_{i}\right\|_{2}^{2}$ $\leq d_{t}c^{2}.$ (6)
When $c=0$, these conditions reduce to exact matching (perfect fit).
The intuition for imposing (4)–(6) is as follows. Suppose the outcomes of the units in the reference and target domains are generated by mechanisms that differ primarily in how they are affected by observed covariates. In this case, if we match on observed covariates and match outcomes over a long period in the reference domain, then we should also be matching on unobserved covariates that affect outcomes in both domains. For instance, in the vaccination-rate application, with the Hispanic and Black sub-populations as the target and reference domains, respectively, factors such as overall attitudes toward vaccines, socioeconomic status, and access to vaccines are unobserved covariates that may similarly affect COVID-19 vaccination rates across both sub-populations. Therefore, if we match on the vaccination rate from one subpopulation for an extended period, as well as on all observed covariates relevant to COVID-19 vaccination, then there is statistical justification that we are also matching on unobserved covariates.
Based on the above argument, a weighted sum of the potential outcomes with weight vector $w$ that satisfies (4)–(6) with a small $c$ will match the potential outcome of the target unit to a good approximation. That is,
@@
Consider the case of $c=0$. In restrictions (4)–(6), we will usually have more equations than unknowns, and, hence, these restrictions cannot be exactly satisfied. Therefore, we instead minimize the mismatch in each restriction. Formally, for $i=1,\dots,J+1$, let $K_{i}:=[\bar{F}_{i}^{\top}~Z_{i}^{\top}~X_{i}^{\top}]^{\top}$. A naive way to obtain the weight vector $w$ is
@@
In the approach above, in line with the original synthetic control method, we require the weights $w$ to be nonnegative and to sum to one. The purpose of this restriction is interpretability and avoiding overfitting. Yet, similar to the standard synthetic control method, we can also prevent overfitting by regularization (Doudchenko and Imbens, 2016).
@@
Here, $\alpha_{i,s}$ is the causal effect on unit $i$ in time period $s$. The vectors $\mu_{i}\in\mathbb{R}^{d_{u}}$ are unit-specific latent factors that affect both domains; $\tilde{\mu}_{i}^{Y}\in\mathbb{R}^{d^{Y}_{\tilde{u}}}$ and $\tilde{\mu}_{i}^{F}\in\mathbb{R}^{d^{F}_{\tilde{u}}}$ are unit-specific latent factors that affect only the target and reference domains, respectively. Loadings are deterministically bounded in $\ell_{2}$-norm: $\sup_{t\leq T}\|\theta_{t}\|_{2}\leq\bar{\theta}$, $\sup_{t\leq T}\|\tilde{\theta}_{t}\|_{2}\leq\bar{\tilde{\theta}}$, $\sup_{s\leq S}\|\vartheta_{s}\|_{2}\leq\bar{\vartheta}$, and $\sup_{s\leq S}\|\tilde{\vartheta}_{s}\|_{2}\leq\bar{\tilde{\vartheta}}$. We also assume the observed-covariate loadings are deterministically bounded: $\sup_{s\leq S}\|\varphi_{s}\|_{2}\leq\bar{\varphi}$ and $\sup_{t\leq T}\|\phi_{t}\|_{2}\leq\bar{\phi}$. The matrix $\sum_{t=1}^{T}\theta_{t}\theta_{t}^{\top}$ has minimum eigenvalue $\lambda_{\min}\geq T\,\underline{\xi}>0$. The shocks satisfy $\mathbb{E}[\epsilon_{i,t}]=\mathbb{E}[\varepsilon_{i,s}]=0$, and $\epsilon_{i,t}$ are i.i.d. sub-Gaussian with variance proxy $\bar{\sigma}^{2}$. The collections $\{\tilde{\mu}^{Y}_{i}\}_{i\leq J+1}$ and $\{\tilde{\mu}^{F}_{i}\}_{i\leq J+1}$ are i.i.d., mean-zero, sub-Gaussian with variance proxy $\tau^{2}$. Let $\mathcal{G}:=\sigma\!\big(\{X_{i},Z_{i},\mu_{i},\tilde{\mu}^{F}_{i}\}_{i\leq J+1},\{\epsilon_{i,t}\}_{i\leq J+1,\,1\leq t\leq T}\big)$ denote the $\sigma$-field generated by all reference-domain objects used to construct the synthetic-control weights $w=(w_{2},\ldots,w_{J+1})$, which are assumed $\mathcal{G}$-measurable, nonnegative, and satisfy $\sum_{i=2}^{J+1}w_{i}=1$. We assume that $\{\varepsilon_{i,s}\}_{i,s}$ and $\{\tilde{\mu}^{Y}_{i}\}_{i}$ are independent of $\mathcal{G}$.
@@
$\big|\mathbb{E}[\hat{\psi}^{sc}]-\psi_{0}\big|\;\leq\;\Bigg(\bar{\varphi}\sqrt{d_{t}}\;+\;\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\;+\;\frac{\bar{\vartheta}\,\bar{\theta}\,\bar{\phi}}{\underline{\xi}}\sqrt{d_{r}}\Bigg)c\;+\;\sqrt{2}\,\frac{\bar{\vartheta}\,\bar{\theta}}{\underline{\xi}}\,\bar{\sigma}\sqrt{\frac{\log(2J)}{T}}\;+\;2\,\frac{\bar{\vartheta}\,\bar{\theta}\,\bar{\tilde{\theta}}}{\underline{\xi}}\,\tau\,\sqrt{\log(2J)}.$
In particular, when $c=0$ (exact matching), the first term vanishes.
Theorem 4 shows that, in the absence of reference-domain–only latent factors ($\tilde{\mu}^{F}_{i}$), the bound on the bias of our proposed estimator vanishes as the number of time periods in the reference domain increases. By contrast, the presence of target-domain–only latent factors ($\tilde{\mu}^{Y}_{i}$) does not prevent convergence. We also see that the bound becomes wider as the number of donor units increases, due to the higher chance of overfitting, i.e., matching by chance. A main ingredient in the proof is a maximal inequality to bound a linear combination of weighted error terms. This step also appears in (Abadie et al., 2010) in the standard synthetic control setting. However, in contrast to that work, we employ a sub-Gaussian approach (see Assumption 5), leading to a clearer and more interpretable derivation without additional moment assumptions on auxiliary variables used in (Abadie et al., 2010). There are also challenges unique to the data fusion setting that do not arise in the standard synthetic control setting: (i) we must handle target-domain–only and reference-domain–only latent factors; and (ii) we must account for two sets of factor loadings for the shared latent factors $\mu_{i}$ across the two domains, which necessitates bounds for both sets. As seen in Theorem 4, the upper bounds on both sets of loadings appear in the bias bound.